OpenStudy (anonymous):

If c dollars are deposited in an account paying r percent annual interest, the amount after x years is a=c(1+r)^x. Approximate the amount in the account after x years when c=$200, r=23%, and x=15 years..any one???

OpenStudy (gw2011):

Substitute into the formula: a=200(1+0.23)^15 a=200(1.23)^15 a=200(22.31) When rounding (1.23)^15 to two decimal places a=$4462 Approximately OR a=200(22.314) When rounding (1.23)^15 to three decimal places a=$4463 Approximately